# Keep the beat going: Automatic drum transcription with momentum

### v1

@@ 3.1 Multiplicative update rule
Proof 3.3.
@@
We will prove that the update rule applied to $\bm{H}_{D}$ does not increase the loss function $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})$. A similar proof shows that the update rule applied to $\bm{W}_{H}$ or $\bm{H}_{H}$ does not increase the loss function.
@@
(5) $\bm{x}_{\star}=\bm{H}_{D}(\cdot,j)\odot\frac{\bm{W}_{D}^{\top}\bm{V}(\cdot,j)}{\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}$
@@
3.2 Projected gradient descent with momentum
Algorithm 2 $\operatorname{OGM}(\bm{W},\bm{H}_{0},\bm{V})$
1:  Initialize $\alpha_{1}=1$, $\bm{Y}_{1}=\bm{H}_{0}$
2:  for $k=1,...,K$ do
4:     $\alpha_{k+1}=\frac{1+\sqrt{4\alpha_{k}^{2}+1}}{2}$
6:  end for
7:  $\bm{H}=\bm{H}_{K}$
8:  return $\bm{H}$
In 2012, Guan and coauthors [6] introduced the “NeNMF” method for the classic NMF problem Eq. 1. Like MUR, the NeNMF method performs an alternating minimization over $\bm{W}$ and $\bm{H}$ factors. As the major innovation, NeNMF solves each subproblem
to high accuracy. To do this, NeNMF applies several iterations of proximal gradient descent with Nesterov momentum, which Guan et al. call the “optimal gradient method” (OGM).
Algorithm 2 provides pseudocode for OGM. The method is based on gradient descent, and it projects each iterate onto the cone of nonnegative-valued matrices by an element-wise application of the rectified linear unit $\operatorname{ReLu}(x)=\max\{x,0\}$. The method incorporates Nesterov momentum using a step size parameter $\alpha_{k}$, which promotes an accelerated convergence rate. Guan et al. established the following guarantee for OGM in [6, Prop. 2.1].
Theorem 3.4 (Convergence of OGM [6]).
$\bm{H}_{\star}\in\operatorname*{arg\,min}_{\bm{H}\geq\bm{0}}\lVert\bm{W}\bm{H}-\bm{V}\rVert_{\rm F}^{2}.$
Theorem 3.4 ensures that OGM solves each subproblem up to an additive error of $\mathcal{O}\big{(}\frac{1}{K^{2}}\big{)}$ after $K$ iterations.
To our knowledge, NeNMF has not previously been applied to the partially fixed NMF problem Eq. 2. However, we can extend NeNMF to the partially fixed NMF problem by using OGM to solve each subproblem in the alternating minimization Eq. 3. Algorithm 3 presents pseudocode for this new NeNMF extension. Because Algorithm 3 is using OGM to solve each subproblem, convergence guarantees follow immediately from Theorem 3.4.
Algorithm 3 NeNMF for partially fixed NMF
3:  for $t=1,...,T$ do
7:  end for
8:  $\bm{H}_{D}=\bm{H}_{D}^{T}$, $\bm{W}_{H}=\bm{W}_{H}^{T}$, $\bm{H}_{H}=\bm{H}_{H}^{T}$.
9:  return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
3.3 Comparison of runtimes and convergence
Updating a single matrix in an iteration of MUR or OGM requires $\mathcal{O}(mn)$ arithmetic operations, and the updates have similar empirical runtimes. Since NeNMF performs $K$ iterations of OGM, the cost of NeNMF is roughly $K$ times higher than the cost of MUR. The results in Section 4 account for this factor-of-$K$ difference by running MUR for $T=100$ iterations versus running NeNMF with $T=10$ outer iterations containing $K=10$ inner OGM iterations.
Obtaining the global minimizer of the partially fixed NMF problem Eq. 2 is challenging since the loss function is nonconvex. Nevertheless, we can hope to solve each convex subproblem in the alternating minimization Eq. 3 to high accuracy. NeNMF is guaranteed to solve each subproblem up to an additive error of $\mathcal{O}(1/K^{2})$ due to Theorem 3.4. In contrast, MUR is not guaranteed to solve each subproblem to any level of accuracy, although it does decrease the overall loss due to Theorem 3.2.
4 Application to automatic drum transcription
This section describes numerical experiments applying the partially fixed NMF for automatic drum transcription. Section 4.1 introduces data sets and Section 4.2 presents results.
4.1 Data sets
“Every Now and Then” is a song from the author’s band, Threadbare [17]. It consists of two guitars, one bass guitar, vocals, and a drum kit, and it is available for streaming at https://open.spotify.com/track/3FoWRPLO9Z0tQiVY9jLrth?si=1e252b92cde04a8e. The sample rate is 44.1 kHz. We manually annotated each onset of the snare drum or bass drum. See Fig. 1 for a labeled photo of the complete drum kit, including the two annotated components.
Figure 1: Photograph of the author’s drum kit with instruments annotated by hand. The snare and bass drums were included in the automatic drum transcription for the song “Every Now and Then”, as described in Section 4.
As a preprocessing procedure, we converted ENST-Drums and “Every Now and Then” into magnitude spectrograms using the short-time Fourier transform with a window length 2048 and hop size 512. This transformation resulted in a time resolution of 32 ms for ENST-Drums and 12 ms for “Every Now and Then”.
4.2 Results
Figure 2: Square Frobenius norm error as a function of runtime for MUR and NeNMF. The horizontal axis shows the number of iterations for MUR and 10×10$\times$10 × the number of outer iterations for NeNMF. The error for ENST-Drums (left) is averaged across 28 tracks, while the error for “Every Now and Then” (right) is for a single track.
We applied the MUR and NeNMF algorithms to solve the partially fixed NMF problem Eq. 2 for ENST-Drums and “Every Now and Then” data using a harmonic rank $r_{H}=5$. In these tests, we independently initialized the entries of $\bm{H}_{D},\bm{W}_{H},$ and $\bm{H}_{H}$ as $\operatorname{Unif}(0,1)$ random variables. Then we applied either MUR for 100 iterations or NeNMF for 10 outer iterations using 10 inner OGM iterations. These experiments led to similar runtimes for the two algorithms.
Figure 2 compares the square Frobenius norm errors achieved using MUR and NeNMF. The errors for ENST-Drums (left panel) are averaged across the 28 tracks, while the errors for “Every Now and Then” (right panel) are for a single track. To promote a fair comparison, the horizontal axis records the iteration parameter $t$ in MUR (Algorithm 1) and it records 10×10$\times$10 × the iteration parameter $t$ in NeNMF (Algorithm 3). The results show that NeNMF leads to a comparable or smaller square Frobenius norm error for either data set and any fixed runtime.
Next, we evaluated the accuracy of the drum hits identified using MUR and NeNMF. Both optimization algorithms produce drum onset matrices consisting of nonnegative entries $\bm{H}_{D}(i,j)$. As a complication, however, there are various ways to determine which entries are large enough to be regarded as drum hits or “onsets”. One standard method used in automatic music transcription [3, 2] identifies $p_{j}$ of the largest entries in each column $\bm{H}_{D}(\cdot,j)$, where $p_{j}$ is the ground-truth number of drum hits. The ground-truth number of drum hits is not likely to be available in real-world use cases. Yet this approach allows for straightforward comparisons without the need to tune evaluation hyperparameters. After identifying the drum onsets using the ground-truth number of onsets, we counted the true positives (TP), false positives (FP), and false negatives (FN) in each row of the onset matrix $\bm{H}_{D}$. Then, we measured accuracy using the $F$-score:
$F=\frac{2{\rm TP}}{2{\rm TP}+{\rm FP}+{\rm FN}}.$
Table 1 reports the F-scores for each data set and optimization method, averaged over all the drum components and tracks for each data set. The results show that NeNMF achieves higher accuracy than MUR for both ENST-Drums and “Every Now and Then”. NeNMF achieves an especially high accuracy of 97.5% for “Every Now and Then”, because the track contains just two annotated drum components and the onset matrix correctly indicates which component is larger at nearly all the onset times.
MUR NeNMF
ENST-Drums 0.599 0.620
“Every Now and Then” 0.854 0.975
Table 1: F-scores evaluated using ground-truth number of offsets. Higher F-scores are marked in bold.
A different way to identify drum hits is based on a signal-adaptive median threshold [19, 10]. This method evaluates the median in a sliding window over each row $\bm{H}_{D}(i,\cdot)$, and it adds a small offset to determine the threshold for a drum hit. For the experiments, we set the window length to 0.1 s, and we set the offset coefficients to 0.05 for the hi-hat, 0.1 for the snare drum, and 0.15 for the bass drum. We considered a detected onset correct if it is within 50 ms of a ground-truth onset. Note that changing the window length and the offset coefficients can lead to large changes in the evaluation scores, and different tracks have better scores for different values.
MUR NeNMF
Hi-hat 0.254 0.283
Snare drum 0.308 0.298
Bass drum 0.290 0.629
Table 2: F-scores for ENST-Drums using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
MUR NeNMF
Snare drum 0.275 0.369
Bass drum 0.444 0.454
Table 3: F-scores for “Every Now and Then” using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
Tables 3 and 3 present the F-scores evaluated using the signal-adaptive median threshold. The scores are averaged over 28 tracks for ENST-Drums. The F-scores are generally higher for NeNMF than for MUR. The highest F-score of 62.9% is achieved using NeNMF to detect the bass drum in ENST-Drums, and it is over 2×2$\times$2 × higher than the corresponding F-score of 29.0% achieved using MUR.
5 Conclusion
In our empirical results, NeNMF is generally more accurate than MUR given a fixed computational budget. Further, NeNMF offers advantageous convergence guarantees. Therefore, we recommend using NeNMF for automatic drum transcription and other partially fixed NMF applications in the future.
Acknowledgments

### v2

@@ 3.1. Multiplicative update rule
Algorithm 1 uses the multiplicative update rule established in 2, preserving the nonnegativity of the iterates and the monotonicity of the loss function.
Algorithm 1 MUR for partially fixed NMF
3:for t=1,,Tt=1,...,T do
7:end for
[…]
9: return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
3.2. Projected gradient descent with momentum
In 2012, Guan and coauthors [10] introduced the “NeNMF” method for the classic NMF problem eq. 1. Like MUR, the NeNMF method is based on an alternating minimization over $\bm{W}$ and $\bm{H}$ factors. As the major innovation, NeNMF solves each subproblem
to high accuracy. To do this, NeNMF applies several iterations of projected gradient descent with Nesterov momentum, which Guan et al. call the “optimal gradient method” (OGM) and which is presented in pseudocode in algorithm 2.
Algorithm 2 $\operatorname{OGM}(\bm{W},\bm{H}_{0},\bm{V})$
1: Initialize $\alpha_{1}=1$, $\bm{Y}_{1}=\bm{H}_{0}$
2:for k=1,,Kk=1,...,K do
4:  $\alpha_{k+1}=\frac{1+\sqrt{4\alpha_{k}^{2}+1}}{2}$
6:end for
[…]
8: return $\bm{H}$
OGM is based on gradient descent, and it projects each iterate onto the cone of nonnegative-valued matrices by an element-wise application of the rectified linear unit $\operatorname{ReLu}(x)=\max\{x,0\}$. The method incorporates Nesterov momentum using a step size parameter $\alpha_{k}$, which promotes an accelerated convergence rate. Guan et al. established the following guarantee for OGM in [10, Prop. 2.1].
Theorem 3 (Convergence of OGM [10]).
$\bm{H}_{\star}\in\operatornamewithlimits{argmin}_{\bm{H}\geq\bm{0}}\lVert\bm{W}\bm{H}-\bm{V}\rVert_{\rm F}^{2}.$
3 ensures that OGM solves each subproblem up to an additive error of $\mathcal{O}\big(\frac{1}{K^{2}}\big)$ after $K$ iterations.
To our knowledge, NeNMF has not previously been applied to the partially fixed NMF problem eq. 3. However, we can extend NeNMF to the partially fixed NMF problem by using OGM to solve each subproblem in the alternating minimization, as presented in algorithm 3. This new NeNMF extension is straightforward, with none of the conceptual or theoretical difficulties as in MUR. Because algorithm 3 is using OGM to solve each subproblem, convergence guarantees follow immediately from 3.
Algorithm 3 NeNMF for partially fixed NMF
3:for t=1,,Tt=1,...,T do
7:end for
[…]
9: return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
3.3. Comparison of theoretical guarantees and runtimes
Obtaining the global minimizer of the partially fixed NMF problem eq. 3 is challenging since the loss function is high-dimensional and nonconvex. Nevertheless, we can hope to solve each convex subproblem in the alternating minimization to high accuracy. NeNMF is guaranteed to solve each subproblem up to an additive error of $\mathcal{O}(1/K^{2})$ due to 3. In contrast, MUR is not guaranteed to solve each subproblem to any level of accuracy, although it does decrease the overall loss due to 2. Therefore, NeNMF has stronger theoretical guarantees.
Updating a single matrix in an iteration of MUR or OGM requires $\mathcal{O}(mnr)$ arithmetic operations where $r=r_{D}+r_{H}$ is the total rank of drum components and harmonic components. Since NeNMF performs $K$ iterations of OGM while MUR performs a single multiplicative update for each matrix, the cost of one iteration of NeNMF is roughly $K$ times higher than the cost of one iteration of MUR. The results in section 4 account for this factor-of-$K$ difference by running MUR for $T=100$ iterations versus running NeNMF with $T=10$ outer iterations containing $K=10$ inner OGM iterations. Therefore, the runtime between algorithms is similar (65 seconds for MUR, 68 seconds for NeNMF).
4. Application to automatic drum transcription
This section describes numerical experiments applying partially fixed NMF to automatic drum transcription. Section 4.1 describes the data sets and section 4.2 presents results.
4.1. Data sets
“Every Now and Then” is a recorded song from the author’s band, Threadbare [24]. It consists of two guitars, one bass guitar, vocals, and a drum kit. It is available for streaming at https://open.spotify.com/track/3FoWRPLO9Z0tQiVY9jLrth?si=1e252b92cde04a8e. The sample rate is 44.1 kHz. We manually annotated each onset of the snare drum or bass drum as a ground-truth reference. See fig. 3 for a labeled photo of the complete drum kit, including the two annotated components.
Figure 3. Photograph of the author’s drum kit with instruments annotated by hand. The snare and bass drums were included in the automatic drum transcription for the song “Every Now and Then”, as described in section 4.
As a preprocessing procedure, we converted ENST-Drums and “Every Now and Then” into magnitude spectrograms using the short-time Fourier transform with a window length 2048 and hop size 512. This transformation resulted in a time resolution of 32 ms for ENST-Drums and 12 ms for “Every Now and Then”. The spectrogram size for “Every Now and Then” was 1025 frequency bins by 13950 time frames. The spectrograms of the tracks in ENST-Drums also included 1025 frequency bins, with varying lengths around 2200 time frames.
4.2. Results
We applied the MUR and NeNMF algorithms to solve the partially fixed NMF problem eq. 3 for ENST-Drums and “Every Now and Then” data using a harmonic rank $r_{H}=5$. In these tests, we independently initialized the entries of $\bm{H}_{D},\bm{W}_{H},$ and $\bm{H}_{H}$ as $\operatorname{Unif}(0,1)$ random variables. Then we applied either MUR for 100 iterations or NeNMF for 10 outer iterations using 10 inner OGM iterations. These experiments led to similar runtimes for the two algorithms. For “Every Now and Then,” the MUR runtime was about 65 seconds, and the NeNMF runtime was about 68 seconds.
Figure 4 compares the square Frobenius norm errors achieved using MUR and NeNMF. The errors for ENST-Drums (left panel) are averaged across the 28 tracks, while the errors for “Every Now and Then” (right panel) are for a single track. To promote a fair comparison, the horizontal axis records the iteration parameter $t$ in MUR (algorithm 1) and it records $10\times$ the iteration parameter $t$ in NeNMF (algorithm 3). Note that the number of inner OGM iterations for NeNMF is fixed at 10 for these experiments. The results show that NeNMF leads to a comparable or smaller square Frobenius norm error for either data set and any fixed runtime.
Figure 4. Square Frobenius norm error as a function of runtime for MUR and NeNMF. The horizontal axis shows the number of iterations for MUR and the number of inner iterations ($10\times$ the number of outer iterations) for NeNMF. Note that the number of inner OGM iterations per outer iteration for NeNMF is fixed at 10. The error for ENST-Drums (left) is averaged across 28 tracks, while the error for “Every Now and Then” (right) is for a single track.
Figure 5. Entry values of drum activation matrix $\bm{H}_{D}$ plotted along time, optimized with NeNMF (top) and MUR (bottom) for the first 30 seconds of “Every Now and Then.” The markers represent ground-truth annotations, and large entry values of $\bm{H}_{D}$ represent detected onsets.
Next, we evaluated the accuracy of the drum hits identified using MUR and NeNMF. Both optimization algorithms produce drum activation matrices consisting of nonnegative entries $\bm{H}_{D}(i,j)$, where a large entry represents a detected onset of drum instrument $i$ at time frame $j$. Figure 5 visually compares the accuracy of the onsets detected by MUR and NeNMF for the first 30 seconds of “Every Now and Then.” The plots show that $\bm{H}_{D}$ optimized with NeNMF more accurately captures the drum onsets.
As a complication, there are various ways to determine which entries of $\bm{H}_{D}$ are large enough to be regarded as drum hits or “onsets”. One standard method used in automatic music transcription [5, 3] identifies $p_{j}$ of the largest entries in each column $\bm{H}_{D}(\cdot,j)$, where $p_{j}$ is the ground-truth number of drum hits. Although the ground-truth number of drum hits is not likely to be available in real-world use cases, this approach allows for straightforward comparisons without the need to tune evaluation hyperparameters. After identifying the drum onsets using the ground-truth number of onsets, we counted the true positives (TP), false positives (FP), and false negatives (FN) in each row of the onset matrix $\bm{H}_{D}$. Then, we measured accuracy using the $F$-score:
$F=\frac{2{\rm TP}}{2{\rm TP}+{\rm FP}+{\rm FN}}.$
Table 1 reports the F-scores calculated using the ground-truth number of offsets for each data set and optimization method, averaged over all the drum components and tracks for each data set. The results show that NeNMF achieves higher accuracy than MUR for both ENST-Drums and “Every Now and Then”. NeNMF achieves an especially high accuracy of 97.5% for “Every Now and Then”, because the track contains just two annotated drum components and the onset matrix correctly indicates which component is larger at nearly all the onset times.
MUR NeNMF
ENST-Drums 0.599 0.620
“Every Now and Then” 0.854 0.975
Table 1. F-scores evaluated using ground-truth number of offsets. Higher F-scores are marked in bold.
A different way to identify drum hits is a signal-adaptive median threshold [26, 16]. This method evaluates the median in a sliding window over each row $\bm{H}_{D}(i,\cdot)$, and it adds a small offset to determine the threshold for a drum hit. For the experiments, we set the window length to 0.1 s, and we set the offset coefficients to 0.05 for the hi-hat, 0.1 for the snare drum, and 0.15 for the bass drum. We considered a detected onset correct if it is within 50 ms of a ground-truth onset. Note that changing the window length and the offset coefficients can lead to large changes in the evaluation scores, and different tracks have better scores for different values.
MUR NeNMF
Hi-hat 0.254 0.283
Snare drum 0.308 0.298
Bass drum 0.290 0.629
Table 2. F-scores for ENST-Drums using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
MUR NeNMF
Snare drum 0.275 0.369
Bass drum 0.444 0.454
Table 3. F-scores for “Every Now and Then” using the signal-adaptive median threshold evaluation. Higher F-scores are marked in bold.
Tables 3 and 3 present the F-scores evaluated using the signal-adaptive median threshold. The scores are averaged over 28 tracks for ENST-Drums. The F-scores are generally higher for NeNMF than for MUR. The highest F-score of 62.9% is achieved using NeNMF to detect the bass drum in ENST-Drums, and this F-score is over $2\times$ higher than the corresponding F-score of 29.0% achieved using MUR.
5. Conclusion
In our empirical results, NeNMF is consistently more accurate than MUR given a fixed computational budget. Further, NeNMF offers advantageous theoretical convergence guarantees. Therefore, we recommend using NeNMF for automatic drum transcription and other partially fixed NMF applications in the future.
As a limitation of this work, the evaluation procedure for the detected drum onsets is sensitive to parameters, and it is unclear which choice works best theoretically or empirically. Further work on a practical and accurate evaluation method would enrich these experiments. As another limitation, this work focuses on two of the many possible optimizers for NMF. Future work comparing other optimization methods or exploring different loss functions like the Kullback-Leibler or Itakura-Saito divergences could be interesting.
In conclusion, this work makes progress toward an accurate and interpretable automatic drum transcription system, first by highlighting partially fixed NMF as an interpretable model choice and second by identifying NeNMF as an efficient optimizer. We acknowledge that the interpretable ADT system in this work does not yet achieve the ideal accuracy level, and indeed a super high-accuracy ADT system (e.g., with $99\%$ accuracy) is not yet available [5, 6, 12, 15, 22, 23, 26, 27, 28]. Indeed, our validation efforts were limited by the amount of accessible drum audio data with accurate annotations, and we manually annotated over 600 drum hits for “Every Now and Then”. Yet, with further research on defining drum “onsets” and further modifications to the loss function, we believe it may be possible to build the partially fixed NMF approach into a fully accurate, interpretable ADT system so that manual annotations become unnecessary in the future.
Acknowledgments
Appendix A Proof of 2
Proof A.1.
@@ Appendix A Proof of 2
This section will prove that the update rule applied to $\bm{H}_{D}$ does not increase the loss function $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})$. A similar proof shows that the update rule applied to $\bm{W}_{H}$ or $\bm{H}_{H}$ does not increase the loss function.
@@
$\bm{x}_{\star}=\bm{H}_{D}(\cdot,j)\odot\frac{\bm{W}_{D}^{\top}\bm{V}(\cdot,j)}{\bm{W}_{D}^{\top}[\bm{W}_{D}\bm{H}_{D}(\cdot,j)+\bm{W}_{H}\bm{H}_{H}(\cdot,j)]}$ (4)